OpenStudy (anonymous):

Factor completely. (0.81y^2-0.0016)

jimthompson5910 (jim_thompson5910):

\[\large 0.81y^2-0.0016\] \[\large (0.9y)^2-(0.04)^2\] \[\large (0.9y-0.04)(0.9y+0.04)\]

jimthompson5910 (jim_thompson5910):

You can represent the answer as \[\large (\frac{9y}{10}-\frac{4}{100})(\frac{9y}{10}+\frac{4}{100})\]